OpenStudy (anonymous):

What is the best way to classify each equation 6x+3(x-4)=8(x-3) A. identity B. contradiction C. neither @shinalcantara

OpenStudy (shinalcantara):

have you remembered what i did last time? now you'll do it this time. 6x +3(x-4) = 8(x-3) distribute 3 to (x-4) and 8 to (x-3)

OpenStudy (shinalcantara):

what will you get?

OpenStudy (anonymous):

b? @shinalcantara

OpenStudy (shinalcantara):

6x +3(x-4) = 8(x-3) 6x + 3x - 12 = 8x - 3 9x - 12 = 8x - 3 are they equal?

OpenStudy (anonymous):

yea @shinalcantara

OpenStudy (shinalcantara):

no. they are not 9x - 12 = 8x - 3 9x is not equal to 8x and -12 is not equal to -3 but there's a solution for we can simplify this as: 9x - 8x = -3 + 12 x = 9 since there's a solution then it's not contradiction then it's C